OpenStudy (sarahc):

Use mathematical induction to prove the statement is true for all positive integers n. 8 + 16 + 24 + ... + 8n = 4n(n + 1)

OpenStudy (freckles):

so you want to show \[\sum_{i=1}^{n}8i=4n(n+1) \] let's see if the base case holds

OpenStudy (freckles):

i will let you test base case

OpenStudy (freckles):

The hardest part is suppose to \[\sum_{i=1}^{k}8i=4k(k+1) \text{ for some integer } k>0\] and then show \[\sum_{i=1}^{k+1}8i=4(k+1)(k+2) \text{ for some integer } k>0\]

OpenStudy (sarahc):

@freckles

OpenStudy (freckles):

looking over

OpenStudy (freckles):

8 + 16 + 24 + . . . + 8k +8(K+1)=4(k+1)((k+1)+1) 8 + 16 + 24 + . . . + 8k +8(K+1)= 4k(k + 1)+8(K+1).....1 what happened between these two lines

OpenStudy (freckles):

or what happened between 8 + 16 + 24 + . . . + 8k = 4k(k + 1) 8 + 16 + 24 + . . . + 8k +8(K+1)=4(k+1)((k+1)+1)

OpenStudy (freckles):

it looks like you added 8(k+1) on one side did you do that to the other side?

OpenStudy (freckles):

\[8+16+24+ \cdots +8k+8(k+1)=4k(k+1)+8(k+1)?\]

OpenStudy (freckles):

8 + 16 + 24 + . . . + 8k +8(K+1)=4(k+1)((k+1)+1) I think I will ignore this line I think ...1 you are just calling that equation 1 in 8 + 16 + 24 + . . . + 8k +8(K+1)= 4k(k + 1)+8(K+1).....1 so that is a good start for this

OpenStudy (freckles):

and it looks like you took it the long way around

OpenStudy (freckles):

this is setup already for factoring by grouping

OpenStudy (freckles):

4k(k+1)+8(k+1) factor the common factor 4(k+1) out

OpenStudy (freckles):

and I get what you did you were just starting what you wanted to show before you started showing it i got confused earlier your proof looks fine

OpenStudy (freckles):

I'm just saying you could reduce the steps by factoring by grouping

OpenStudy (sarahc):

Ah, yes thats true. I understand, thank you

OpenStudy (freckles):

np